# Inertial antiferromagnetic resonance driven by spin-orbit torques

### v1

@@ Abstract
It is widely accepted that the handedness of a resonant mode is an intrinsic property. We show that, by tailoring the polarization and handedness of alternating spin-orbit torques used as the driving force, the polarization state and handedness of inertial resonant modes in an antiferromagnet can be actively controlled. In contrast to ferromagnets, whose resonant-mode polarization is essentially fixed, antiferromagnetic inertial modes can continuously evolve from elliptic through circular to linear polarization as the driving polarization is varied. We further identify an inertia-dependent critical degree of driving polarization at which the mode becomes linearly polarized while its handedness reverses.
@@
Spin inertia [1] in magnetically ordered materials can be induced by various microscopic mechanisms, as demonstrated by recent studies. First-principles spin dynamics have shown that both damping and inertia emerge consistently from the electronic structure, emphasizing the role of spin-orbit coupling in dissipative and inertial spin dynamics [2, 3]. A complementary approach based on spin-bath coupling reveals that a non-Markovian memory kernel yields an effective inertial term in the spin equation of motion [4, 5]. Additionally, nonadiabatic interactions with environmental degrees of freedom can produce inertia-like corrections [6]. A classical interpretation models the spin as a circular current loop, leading to inertial terms from the underlying mechanical dynamics [7, 8].
@@
Spin inertia gives rise to magnetic nutation [9, 10, 6, 11, 12], which manifests as nutation resonance [13, 14, 15, 16]. Furthermore, inertia modifies the self-oscillatory dynamics of magnetization in both antiferromagnets (AFMs) [17, 18] and ferromagnets (FMs) [19]. The characteristic influence of spin inertia on these uniform magnetic dynamics can serve as an additional experimental probe for its detection.
@@
Apart from experimental observations of nutation [13, 14, 15, 16], extensive efforts have been invested in the theoretic research of nutational resonance [20, 21, 22, 26, 27, 29, 28, 23, 24, 25]. It was found that the inertial magnetization dynamics are described by the inertial Landau-Lifshitz-Gilbert (ILLG) equation, which predicts a nutation resonance in addition to conventional ferromagnetic resonance. Olive et al. [20, 21] numerically demonstrated a secondary high-frequency peak whose amplitude and position depend on the inertial relaxation time, the damping, and the applied field. Subsequent analytical studies quantified shifts of resonant frequencies and linewidths of both precession and nutation, revealing a non-trivial dependence on damping and magnetic anisotropy [22, 24, 23, 25]. Another analytical study predicts that these effects are enhanced in AFMs due to exchange interaction [26]. Remarkably, when driven by a circularly polarized magnetic field, FMs exhibit a single nutational peak and a single precessional peak, whereas AFMs and ferrimagnets present these peaks in pairs. In the latter two systems, inter- and intra-sublattice couplings further tune these resonances, indicating distinct mechanisms in multi-sublattice systems [27, 28]. Finally, spin pumping at terahertz nutation frequencies produces a spin current opposite in direction to that generated by precessional modes, providing an unambiguous experimental signature of nutation [29].
@@
We note that most studies on inertial resonances have primarily focused the resonant frequency, determined by the intrinsic magnetic properties. However, the polarization of the resonant mode, which depends on both the magnetic material and the external drive, remains comparatively unexplored, despite offering an extra degree of freedom – in addition to amplitude and phase – for information encoding [30]. Experimentally, resonance is usually driven by circularly or linearly polarized magnetic fields. Accordingly, the nutational resonances in the current theoretical studies were assumed to be excited by circularly [22, 26, 27, 29, 28, 25] or linearly polarized [20, 21] oscillating fields.
An alternative utilises current-induced spin-torque resonance, realised via spin-transfer torques in spin-valve structures [31, 32, 33, 34, 35, 36] and spin-orbit torques (SOTs) in heavy-metal/FM bilayers [37, 38, 39, 40], both of which are efficient and scalable in spintronic devices. In both schemes an alternating charge current generates an oscillating spin current that exerts a spin torque on the magnetization. Whereas spin-transfer torque employs a current perpendicular to the film, spin-orbit torque relies on an in-plane current, enabling finer control of the polarization (linear, elliptical or circular) and handedness (left- or right-handed) of the driving torque and, consequently, of the resonant mode.
Motivated by the above considerations, we present the investigation of inertial resonant modes in AFMs driven by alternating spin-orbit torques, with an emphasis on their polarization and handedness.
@@
As illustrated in Fig. 1, we study an inertial AFM film attached to a heavy metal (HM) with strong spin-orbit coupling, driven by oscillating in-plane currents. The bipartite AFM dynamics, that includes magnetic inertia, Gilbert damping, and SOTs, are captured phenomenologically by two exchange-coupled ILLG equations. For sublattice $k$, the ILLG equation reads:
@@
where $\mathbf{m}_{k}$ is the unit magnetization vector of sublattice $k$ ($k=1,2$), $\alpha$ is the Gilbert damping constant, and $\eta$ is the inertial relaxation time, typically ranging from femtoseconds to picoseconds, as predicted by ab initio calculations [41] and confirmed experimentally [13, 14, 15]. The magnetic energy, arising from the magnetocrystalline anisotropy and the intersublattice exchange interaction, is expressed in frequency units as
@@
where $\omega_{E(K)}=\gamma_{0}H_{E(K)}$, with $H_{E}$ and $H_{K}$ being the exchange field and the anisotropy field, respectively. $\gamma_{0}=g\mu_{0}\mu_{B}/\hbar$ is the gyromagnetic ratio with $g$ being the Landé $g$-factor, $\mu_{0}$ the vacuum susceptibility, $\mu_{B}$ the Bohr magneton, and $\hbar$ the reduced Plank constant. The uniaxial anisotropy axis is oriented perpendicular to the film plane.
The last term in Eq. (1) represents the SOTs, originating from spin transfer between the local magnetization and spin currents generated in the adjacent HM layer. Including both dampinglike and fieldlike components [42], the SOTs are given by
$\mathbf{\tau}_{k}=-\rho\left\{\mathbf{m}_{k}\!\times\!\left[\mathbf{m}_{k}\!\times\!\left(\mathbf{e}_{z}\!\times\!\mathbf{j}_{e}\right)\right]\!+\!\beta\mathbf{m}_{k}\!\times\!\left(\mathbf{e}_{z}\!\times\!\mathbf{j}_{e}\right)\right\},$ (3)
where $\beta$ is the ratio of fieldlike to dampinglike SOT. The prefactor $\rho=\xi\mu_{B}/(eM_{s}d)$ characterizes the SOT strength per unit current density, with $M_{s}$ being the saturation magnetization of sublattice, $d$ the thickness of the AFM layer, and $e$ the elementary charge. The SOT efficiency is defined as $\xi=T_{\mathrm{int}}\theta_{\mathrm{sh}}$ [43, 44], where $\theta_{\mathrm{sh}}$ is the spin Hall angle [42] and $T_{\mathrm{int}}$ the spin transparency of the HM/AFM interface [45]. The in-plane electric current density $\mathbf{j}_{e}$ is expressed in components as
@@
Due to the uniaxial anisotropy and intersublattice exchange [Eq. (2)], the static equilibrium magnetizations of the two sublattices align along the ±$\pm z$ directions, respectively. For small deviations from equilibrium, the $z$-components of the magnetizations remain unchanged to first order, such that $m_{1}^{z}\approx 1$ and $m_{2}^{z}\approx-1$. To describe the magnetic response of the AFM in the linear regime, we adopt the ansatz $\mathbf{m}_{k}=(-1)^{k+1}\mathbf{e}_{z}+m_{k}^{x}(t)\mathbf{e}_{x}+m_{k}^{y}(t)\mathbf{e}_{y}$, where $m_{k}^{x}(t)$, $m_{k}^{y}(t)$ are complex-valued functions encoding the amplitude and phase of the response to SOTs. Inserting the ansatz into Eq. (1), assuming a harmonic time dependence of $m_{k}^{x}(t)$, $m_{k}^{y}(t)$ (i.e., replacing $d/dt$ by $-i\omega$), and keeping only terms linear in $m_{k}^{x}(t)$, $m_{k}^{y}(t)$, $j_{x}$ and $j_{y}$, one obtains a four-dimensional linear inhomogeneous system of ordinary differential equations that describes damped magnetic oscillations driven by periodic forces. Solving this system yields the steady-state linear response modes:
$\left(\begin{array}[]{c}m_{1}^{x}(t)\\ m_{1}^{y}(t)\\ m_{2}^{x}(t)\\ m_{2}^{y}(t)\end{array}\right)=\chi\left(\begin{array}[]{c}j_{x}(t)\\ j_{y}(t)\\ j_{x}(t)\\ j_{y}(t)\end{array}\right).$ (5)
Here, in analogy with the Polder susceptibility tensor [46] describing the linear response to an alternating magnetic field, we define a susceptibility tensor characterizing the linear response to alternating SOTs, which reads
$\chi=\frac{\rho}{\Delta}\left(\begin{array}[]{cccc}\chi_{1}&\chi_{2}&-\omega_{E}&\beta\omega_{E}\\ -\chi_{2}&\chi_{1}&-\beta\omega_{E}&-\omega_{E}\\ \omega_{E}&\beta\omega_{E}&-\chi_{1}&\chi_{2}\\ -\beta\omega_{E}&\omega_{E}&-\chi_{2}&-\chi_{1}\end{array}\right),$ (6)
where $\chi_{1}=(\eta\omega^{2}-\omega_{E}-\omega_{K})+i(\alpha-\beta)\omega$, $\chi_{2}=\beta(\eta\omega^{2}-\omega_{E}-\omega_{K})+i(1+\alpha\beta)\omega$, and
@@
$\mathcal{A}^{x(y)}_{1,2}=\rho\sqrt{\frac{P_{1}J^{2}_{x(y)}+P_{2}J^{2}_{y(x)}\pm P_{3}J_{x}J_{y}}{\widetilde{\Delta}}},$ (11)
@@
Figure 2: (color online). Resonant spectra for various inertia parameters $\eta$. Peaks enclosed in the circles (ellipses) correspond to precessional (nutational) resonances. Both axes are shown on a logarithmic scale; low-frequency parts are omitted for clarity. Curves are evaluated from Eq. (11) using $\omega_{E}=9.25$ THz, $\omega_{K}=0.14$ THz, $\alpha=0.01$, $\beta=0.02$, $J_{x}=1$ GA///m2, $J_{y}=2$ GA///m2, and $\rho\approx 0.13$ Hz///(A///m2). The values of $\omega_{E}$, $\omega_{K}$, and $\rho$ are derived from the magnetic parameters of MnF2 [48] for a film thickness $d=3$nm and SOT efficiency $\xi=0.32$.
@@
where $\operatorname{\tan}^{-1}(x,y)$ denotes the two-argument arctangent that returns the angle of the vector $(x,y)$, correctly resolving the quadrant [47]. In Eqs. (13) and (14), the coefficients before $J_{x,y}$, relying on $\omega$, are given by $Q_{1}=(1+\alpha\beta)\Omega_{1}\omega^{2}-[\alpha(\alpha-\beta)\omega^{2}+\Omega_{1}^{2}]\Omega_{2}$, $Q_{2}=\omega\{(1+\alpha^{2})(1+\alpha\beta)\omega^{2}+[\alpha\beta\Omega_{2}-\Omega_{1}]\Omega_{2}\}$, $Q_{3}=\omega\{(1+\alpha^{2})(\alpha-\beta)\omega^{2}+[\alpha\Omega_{1}+\beta\Omega_{2}]\Omega_{1}\}$, and $Q_{4}=(\alpha-\beta)\Omega_{2}\omega^{2}+[\alpha(1+\alpha\beta)\omega^{2}+\beta\Omega_{2}^{2}]\Omega_{1}$. Eqs. (13) and (14) show that the phase lag of the stationary forced oscillations relative to the driving SOTs depends on the amplitude ratio $J_{x}/J_{y}$ and on the driving frequency $\omega$.
@@
where $X_{1,2}=\pm(Q_{1}Q_{4}-Q_{2}Q_{3})(J_{x}^{2}-J_{y}^{2})$, and $Y_{1,2}=\pm(Q_{1}Q_{2}+Q_{3}Q_{4})(J_{x}^{2}+J_{y}^{2})+(Q_{1}^{2}+Q_{2}^{2}+Q_{3}^{2}+Q_{4}^{2})J_{x}J_{y}$.
When the SOT frequency matches the intrinsic eigenfrequencies of inertial AFMs, the amplitudes of magnetization oscillations exhibit resonant enhancement, as one can see in the resonant spectra. Fig. 2 shows the precessional and nutational spectra for different inertial relaxation times $\eta$, considering the general case $J_{x}\neq J_{y}$. The left (right) column displays spectra at negative (positive) frequencies. It is important to note that the panels (a) and (b), (c) and (d), (e) and (f), and (g) and (h) reveals an asymmetry of the spectra with respect to frequency inversion for a given $\eta$. Additionally, comparing panels (a) and (f), (b) and (e), (c) and (h), and (d) and (g), one observes that the positive-frequency spectrum of $m_{1}^{x,y}$ mirrors the negative-frequency spectrum of $m_{2}^{x,y}$, and vice versa. This can be understood from symmetry arguments: exchanging the sublattices ($\mathcal{A}_{1}^{x,y}\leftrightarrow\mathcal{A}_{2}^{x,y}$) is equivalent to time reversal ($\omega\leftrightarrow-\omega$). Therefore, the spectra are symmetric under the combined transformation $1\leftrightarrow 2$ and $\omega\leftrightarrow-\omega$.
Figure 3: (color online). (a) Resonant frequencies as a function of $\eta$. Curves are computed from Eq. (17), while markers denote peak positions extracted from the resonant spectra. A logarithmic scale is used on the frequency axis, and low-frequency regions are omitted for clarity. (b) and (c) show the dependence of peak heights on $\eta$ for precession and nutation, respectively. Curves are computed from Eqs. (20) and (19), and markers indicate values extracted from the spectra. The parameters used are the same as in Fig. 2.
Fig. 2 also shows that the precessional resonances are stronger (higher peaks) and occur at lower frequencies compared to nutational ones. However, both resonances lie in the terahertz range, due to the exchange enhancement characteristic of AFMs.
IV Frequency and magnitude of resonances

### v2

# Chirality and polarization of inertial antiferromagnetic resonances driven by spin-orbit torques

@@ Abstract
It is widely accepted that the handedness of a resonant mode is an intrinsic property. We show that, by tailoring the polarization and handedness of alternating spin-orbit torques used as the driving force, the polarization state and handedness of inertial resonant modes in an antiferromagnet (AFM) can be actively controlled. In contrast with ferromagnets, whose resonant-mode polarization is essentially fixed, AFM inertial modes can continuously evolve from elliptic through circular to linear polarization as the driving polarization is varied. We further identify an inertia-dependent critical degree of driving polarization at which the mode becomes linearly polarized while its handedness reverses.
@@
Spin inertia [RMondal] in magnetically ordered materials can be induced by various microscopic mechanisms, as demonstrated in recent studies. First-principles spin dynamics have shown that both damping and inertia emerge consistently from the electronic structure, emphasizing the role of spin-orbit coupling in dissipative and inertial spin dynamics [SBhattacharjee, RMondal_JPCM]. A complementary approach based on spin-bath coupling reveals that a non-Markovian memory kernel yields an effective inertial term in the spin equation of motion [MAQuarenta, JAnders]. Additionally, nonadiabatic interactions with environmental degrees of freedom can produce inertia-like corrections [TKikuchi]. A classical interpretation models the spin as a circular current loop, leading to inertial terms from the underlying mechanical dynamics [SGiordano, JEWegrowe].
@@
Spin inertia gives rise to magnetic nutation [MCCiornei, DBottcher, TKikuchi, SVTitov_PRB103, SVTitov], which manifests as nutation resonance [YLi, KNeeraj, VUnikandanunni, ADe]. Furthermore, inertia modifies the self-oscillatory dynamics of magnetization in both antiferromagnets (AFMs) [PBHe_PRB108, PBHe_PRB110] and ferromagnets (FMs) [RRodriguez]. The characteristic influence of spin inertia on these uniform magnetic dynamics can serve as an additional experimental probe for its detection.
@@
Apart from experimental observations of nutation [YLi, KNeeraj, VUnikandanunni, ADe], extensive efforts have been invested in the theoretic research of nutational resonance [EOlive_APL, EOlive_JAP, MCherkasskii_PRB102, RMondal_PRB103, RMondal_PRB104, RMondal_PRB104_21, MondalR_JPCM33, SVTitov_JAP, MCherkasskii_PRB106, SGhosh]. It was found that the inertial magnetization dynamics are described by the inertial Landau-Lifshitz-Gilbert (ILLG) equation, which predicts a nutation resonance in addition to conventional FM resonance. Olive et al. [EOlive_APL, EOlive_JAP] numerically demonstrated a secondary high-frequency peak whose amplitude and position depend on the inertial relaxation time, the damping, and the applied field. In subsequent analytical studies shifts of resonant frequencies and linewidths of both precession and nutation have been quantified, revealing a nontrivial dependence on damping and magnetic anisotropy [MCherkasskii_PRB102, MCherkasskii_PRB106, SVTitov_JAP, SGhosh]. Authors of another analytical study have predicted that these effects are enhanced in AFMs due to exchange interaction [RMondal_PRB103]. Remarkably, when driven by a circularly polarized magnetic field, FMs exhibit a single nutational peak and a single precessional peak, whereas AFMs and ferrimagnets present these peaks in pairs. In the latter two systems, inter- and intrasublattice couplings further tune these resonances, indicating distinct mechanisms in multisublattice systems [RMondal_PRB104, MondalR_JPCM33]. Finally, spin pumping at terahertz nutation frequencies produces a spin current opposite in direction to that generated by precessional modes, providing an unambiguous experimental signature of nutation [RMondal_PRB104_21].
@@
We note that authors of most studies on inertial resonances have primarily focused the resonant frequency, determined by the intrinsic magnetic properties. However, the polarization of the resonant mode, which depends on both the magnetic material and the external drive, remains comparatively unexplored, despite offering an extra degree of freedom – in addition to amplitude and phase – for information encoding [RCheng, TYu, WYu, CJia]. Experimentally, resonance is usually driven by circularly or linearly polarized magnetic fields. Accordingly, the nutational resonances in the current theoretical studies were assumed to be excited by circularly [MCherkasskii_PRB102, RMondal_PRB103, RMondal_PRB104, RMondal_PRB104_21, MondalR_JPCM33, SGhosh] or linearly polarized [EOlive_APL, EOlive_JAP] oscillating fields.
An alternative utilizes current-induced spin-torque resonance, realized via spin-transfer torques in spin-valve structures [AATulapurkar, JCSankey, HXi, JNKupferschmidt, AAKovalev, PBHe] and spin-orbit torques (SOTs) in heavy-metal/FM bilayers [LLiu, KKondou, VSluka, CSun], both of which are efficient and scalable in spintronic devices. In both schemes an alternating charge current generates an oscillating spin current that exerts a spin torque on the magnetization. Particularly, in SOT-based devices the in-plane current is more easily tuned than the perpendicular current, enabling finer control of the alternating driving torque. In addition, the role of the inertial effects in mode polarization has received limited attention in the literature. Authors of prior studies on spin dynamics driven by SOTs in inertial magnetic materials have implicitly assumed that the handedness and polarization are essentially fixed by the material. A systematic investigation of the polarization of resonant modes and its tuning by driving forces is, to the best of our knowledge, lacking in the literature.
Motivated by the above considerations, we present the investigation of inertial resonant modes in AFMs driven by alternating SOTs. We show that SOTs generated by two orthogonal in-plane currents tune the polarization state (linear, elliptical or circular) and handedness (left or righthanded) of both precessional and nutational AFM resonances. We find that the handedness reverses at critical current ratios that depend explicitly on the inertial relaxation time. These inertia-dependent switching points provide a practical route to determine spin inertia across a broad range of materials. We use chirality and handedness interchangeably, following common usage in the literature.
@@
As illustrated in Fig. 1, we study an inertial AFM film attached to a heavy metal (HM) with strong spin-orbit coupling, driven by oscillating in-plane currents. The bipartite AFM dynamics are captured phenomenologically by two exchange-coupled ILLG equations, which include spin inertia, Gilbert damping, and SOTs. For sublattice $k$, the ILLG equation reads:
@@
where $\mathbf{m}_{k}$ is the unit magnetization vector of sublattice $k$ ($k=1,2$), $\alpha$ is the Gilbert damping constant, and $\eta$ is the inertial relaxation time, typically ranging from femtoseconds to picoseconds, as predicted by ab initio calculations [DThonig] and confirmed experimentally [YLi, KNeeraj, VUnikandanunni].
The magnetic energy, arising from the magnetocrystalline anisotropy and the intersublattice exchange interaction, is expressed in frequency units as
@@
where $\omega_{E(K)}=\gamma_{0}H_{E(K)}$, with $H_{E}$ and $H_{K}$ being the exchange field and the anisotropy field, respectively. Here, $\gamma_{0}=g\mu_{0}\mu_{B}/\hbar$ is the gyromagnetic ratio with $g$ being the Landé $g$ factor, $\mu_{0}$ the vacuum susceptibility, $\mu_{B}$ the Bohr magneton, and $\hbar$ the reduced Plank constant. The anisotropy easy axis is oriented perpendicular to the film plane.
The last term in Eq. (1) represents the SOTs, originating from spin transfer between the local magnetization and spin currents generated in the adjacent HM layer. Including both dampinglike and fieldlike components [AManchon], the SOTs are given by
$\mathbf{\boldsymbol{\tau}}_{k}=-\rho\left\{\mathbf{m}_{k}\!\times\!\left[\mathbf{m}_{k}\!\times\!\left(\mathbf{e}_{z}\!\times\!\mathbf{j}_{e}\right)\right]\!+\!\beta\mathbf{m}_{k}\!\times\!\left(\mathbf{e}_{z}\!\times\!\mathbf{j}_{e}\right)\right\},$ (3)
where $\beta$ is the ratio of fieldlike to dampinglike SOT.
The prefactor $\rho=\xi\mu_{B}/(eM_{s}d)$ sets the overall SOT strength. Here $M_{s}$ is the saturation magnetization of sublattice, $d$ denotes the thickness of the AFM layer, and $e$ is the elementary charge. The SOT efficiency is defined as $\xi=T_{\mathrm{int}}\theta_{\mathrm{sh}}$ [LZhu, CFPai], where $\theta_{\mathrm{sh}}$ is the spin Hall angle [AManchon] and $T_{\mathrm{int}}$ the spin transparency [ZLWang] of the HM/AFM interface. The dampinglike SOT is nonconservative. It represents the antidamping torque that is even under $\mathbf{m}_{k}\to-\mathbf{m}_{k}$. It can be written as a torque from an effective field $\mathbf{H}_{\mathbf{DL}}\propto\mathbf{m}_{k}\times\mathbf{p},$ where $\mathbf{p}\equiv\mathbf{e}_{z}\times\mathbf{j}_{e}$, thereby modifying the effective Gilbert damping in linear response. The fieldlike SOT is precessional conservative torque that is odd under $\mathbf{m}_{k}\to-\mathbf{m}_{k}$. It induces precession about an effective field $\mathbf{H}_{\mathrm{FL}}\propto\mathbf{p},$where $\mathbf{p}$ is the spin-polarization direction set by the spin Hall effect in the HM.
The in-plane electric current density $\mathbf{j}_{e}$ is expressed in components as
@@
Each term in the ILLG equation encodes a distinct physical mechanism. The precessional term, $\mathbf{m}_{k}\times d\mathcal{E}/d\mathbf{m}_{k}$ drives conservative motion in the effective field $\mathbf{H}_{{\rm eff}}=d\mathcal{E}/d\mathbf{m}_{k}$ set by the magnetic potential energy. The Gilbert term represents dissipation due to coupling of the magnetic moments to a thermal bath. The spin-inertial term accounts for the dynamics related to the kinetic energy. The spin–orbit terms captures the external torques exerted by the HM layer on the AFM film. Although each contribution is well understood, their competition produces the nontrivial dynamics analyzed in Secs. III-V.
One of the issues in inertial spin dynamics is the applicable time scale, set by the inertial relaxation time. Reported values span orders of magnitude: estimates near $1$ fs were proposed in Ref. [SBhattacharjee] and inferred from FM-resonance precession data in Ref. [YLi]. First-principles calculations [DThonig] gave values $\eta\approx 1-100$ fs. Ab initio simulations [Bouaziz2019] of the dynamical magnetic susceptibility indicated $\eta\sim 10-100$ fs. Time-resolved magneto-optical detection of resonant nutation [KNeeraj] yielded $\eta\sim 100$ fs. Measurements on cobalt films [VUnikandanunni] reported $\eta/\alpha\approx 750$ fs. Given the wide range of reported values for the inertial relaxation time, we present results as a function of $\eta.$
@@
Due to the uniaxial anisotropy and intersublattice exchange [Eq. (2)], the static equilibrium magnetizations of the two sublattices align along the $\pm z$ directions, respectively. For small deviations from equilibrium, the $z$-components of the magnetizations remain unchanged to first order, such that $m_{1}^{z}\approx 1$ and $m_{2}^{z}\approx-1$. To describe the magnetic response of the AFM in the linear regime, we adopt the ansatz $\mathbf{m}_{k}=(-1)^{k+1}\mathbf{e}_{z}+m_{k}^{x}(t)\mathbf{e}_{x}+m_{k}^{y}(t)\mathbf{e}_{y}$, where $m_{k}^{x}(t)$, $m_{k}^{y}(t)$ are complex-valued functions encoding the amplitude and phase of the response to SOTs. Inserting the ansatz into Eq. (1), assuming a harmonic time dependence of $m_{k}^{x}(t)$, $m_{k}^{y}(t)$ (i.e., replacing $d/dt$ by $i\omega$) and keeping only terms linear in $m_{k}^{x}(t)$, $m_{k}^{y}(t)$, $j_{x}$ and $j_{y}$, one obtains a four-dimensional linear inhomogeneous system of ordinary differential equations that describes damped magnetic oscillations driven by periodic forces. As detailed in Appendix A, solving this system yields the steady-state linear response modes:
$\left[\begin{array}[]{c}m_{1}^{x}(t)\\ m_{1}^{y}(t)\\ m_{2}^{x}(t)\\ m_{2}^{y}(t)\end{array}\right]=\frac{\rho}{\Delta}\left(\begin{array}[]{cc}\chi_{1}&\chi_{2}\\ -\chi_{2}&\chi_{1}\\ -\chi_{1}&\chi_{2}\\ -\chi_{2}&-\chi_{1}\end{array}\right)\left[\begin{array}[]{c}j_{x}(t)\\ j_{y}(t)\end{array}\right].$ (5)
where $\chi_{1}=(\eta\omega^{2}-2\omega_{E}-\omega_{K})+i(\alpha-\beta)\omega$, $\chi_{2}=\beta(\eta\omega^{2}-\omega_{K})+i(1+\alpha\beta)\omega$, and
@@
In Eq. (5), a susceptibility tensor is defined to characterize the linear response to alternating SOTs, in analogy with the Polder susceptibility tensor [DDStancil] describing the linear response to an alternating magnetic field.
@@
$\mathcal{A}_{1,2}^{x(y)}=\rho\sqrt{\frac{P_{1}J_{x(y)}^{2}+P_{2}J_{y(x)}^{2}\pm P_{3}J_{x}J_{y}}{\widetilde{\Delta}}}$, (10)
@@
Figure 2: (color online). AFM spectra without spin inertia exhibit the single precessional resonance (dashed lines, $\eta=0$). Finite spin inertia $\eta>0$ shifts the precessional AFM resonance and induces the additional nutational resonances (solid lines). Peaks labeled with red (purple) points correspond to precessional (nutational) resonances. Both axes are shown on a logarithmic scale; low-frequency parts are omitted for clarity. Curves are evaluated from Eq. (10) using $\omega_{E}=9.25$ THz, $\omega_{K}=0.14$ THz, $\alpha=0.01$, $\beta=0.02$, $J_{x}=1$ GA//m2, $J_{y}=2$ GA//m2, and $\rho\approx 0.13$ Hz//(A//m2). The values of $\omega_{E}$, $\omega_{K}$, and $\rho$ are derived from the magnetic parameters of MnF2 [PVaidya] for a film thickness d=3d=3nm and SOT efficiency $\xi=0.32$.
@@
where $\tan^{-1}\left(x,y\right)$ denotes the two-argument arctangent that returns the angle of the vector $(x,y)$, correctly resolving the quadrant [arctan]. In Eqs. (12) and (13), the coefficients before $J_{x,y}$, relying on $\omega$, are given by $Q_{1}=(1+\alpha\beta)\Omega_{1}\omega^{2}-[\alpha(\alpha-\beta)\omega^{2}+\Omega_{1}^{2}]\Omega_{2}$, $Q_{2}=\omega\{(1+\alpha^{2})(1+\alpha\beta)\omega^{2}+[\alpha\beta\Omega_{2}-\Omega_{1}]\Omega_{2}\}$, $Q_{3}=\omega\{(1+\alpha^{2})(\alpha-\beta)\omega^{2}+[\alpha\Omega_{1}+\beta\Omega_{2}]\Omega_{1}\}$, and $Q_{4}=(\alpha-\beta)\Omega_{2}\omega^{2}+[\alpha(1+\alpha\beta)\omega^{2}+\beta\Omega_{2}^{2}]\Omega_{1}$. Eqs. (12) and (13) show that the phase lag of the stationary forced oscillations relative to the driving SOTs depends on the amplitude ratio $J_{x}/J_{y}$ and on the driving frequency $\omega$.
@@
where $X_{1,2}=\pm(Q_{1}Q_{4}-Q_{2}Q_{3})(J_{x}^{2}-J_{y}^{2})$, and $Y_{1,2}=\pm(Q_{1}Q_{2}+Q_{3}Q_{4})(J_{x}^{2}+J_{y}^{2})+(Q_{1}^{2}+Q_{2}^{2}+Q_{3}^{2}+Q_{4}^{2})J_{x}J_{y}$. Note that when flipping the sign of $\omega$, $Q_{1}$ and $Q_{4}$ remain unchanged and $Q_{2}$ and $Q_{3}$ change the sign. Consequently, $X_{1,2}$ do not change and $Y_{1}$ ($Y_{2}$) becomes $Y_{2}$ ($Y_{1}$).
The system shows two SOT–driven resonances: a low-frequency precession and a high-frequency nutation created by spin inertia (Fig. 2). Inertia enters via the $\eta\omega^{2}$ terms in $\Omega_{1,2}$, shifting the spectrum and producing the nutation peak. One can see that the precessional resonances are stronger (higher peaks) and occur at lower frequencies than nutational ones. However, both resonances lie in the terahertz range, due to the exchange enhancement characteristic of AFMs. Figure 2 shows the precessional and nutational spectra for different inertial relaxation times $\eta$, considering the general case $J_{x}\neq J_{y}$. The left (right) column displays spectra at negative (positive) frequencies. It is important to note that panels (a) and (b), (c) and (d), (e) and (f), and (g) and (h) reveal an asymmetry of the spectra with respect to frequency inversion for a given sublattice. Additionally, comparing panels (a) and (f), (b) and (e), (c) and (h), and (d) and (g), one observes that the positive-frequency spectrum of $m_{1}^{x,y}$ mirrors the negative-frequency spectrum of $m_{2}^{x,y}$ and vice versa. This can be understood from symmetry arguments: Exchanging the sublattices ($\mathcal{A}_{1}^{x,y}\leftrightarrow\mathcal{A}_{2}^{x,y}$) is equivalent to time reversal ($\omega\leftrightarrow-\omega$). Therefore, the spectra are symmetric under the combined transformation $1\leftrightarrow 2$ and $\omega\leftrightarrow-\omega$.
Figure 3: (color online). (a) Resonant frequencies as a function of $\eta$. Curves are computed from Eq. (16). A logarithmic scale is used on the frequency axis, and low-frequency regions are omitted for clarity. (b) and (c) show the dependence of peak heights on $\eta$ for precession and nutation, respectively. Curves are computed from Eqs. (19) and (18). The parameters used are the same as in Fig. 2.
IV Resonant peak and frequency